# CASA: Class-Agnostic Shared Attributes in Vision-Language Models for Efficient Incremental Object Detection

### v1

@@ Abstract
Incremental object detection (IOD) is challenged by background shift, where background categories in sequential data may include previously learned or future classes. Inspired by the vision-language foundation models such as CLIP, these models capture shared attributes from extensive image-text paired data during pre-training. We propose a novel method utilizing attributes in vision-language foundation models for incremental object detection. Our method constructs a Class-Agnostic Shared Attribute base (CASA) to capture common semantic information among incremental classes. Specifically, we utilize large language models to generate candidate textual attributes and select the most relevant ones based on current training data, recording their significance in an attribute assignment matrix. For subsequent tasks, we freeze the retained attributes and continue selecting from the remaining candidates while updating the attribute assignment matrix accordingly. Furthermore, we employ OWL-ViT as our baseline, preserving the original parameters of the pre-trained foundation model. Our method adds only 0.7% to parameter storage through parameter-efficient fine-tuning to significantly enhance the scalability and adaptability of IOD. Extensive two-phase and multi-phase experiments on the COCO dataset demonstrate the state-of-the-art performance of our proposed method.
1 Introduction
@@
State-of-the-art object detection models [10, 36, 35, 13, 3, 26] have achieved remarkable success in recognizing and localizing objects within a fixed set of categories. These models operate under a critical limitation: they assume a static training environment where all object categories are known and labelled beforehand. In real-world scenarios, as new object categories emerge over time, the need for continuous updates to these models becomes crucial—a process known as incremental object detection (IOD) [37, 9, 24]. However, IOD also suffers significant challenges, particularly in managing background drift [4]. Background drift occurs when objects belonging to previous or future tasks in IOD are not annotated in the current task and are instead assigned to the background class. This can cause serious misclassification issues [5, 24], as the model may later confuse these background objects with newly introduced categories or fail to recognize them altogether. This highlights a key challenge in incremental learning: how to maintain the ability to generalize across tasks without comprehensive annotations.
@@
In traditional object detection [11, 22, 53], accurate classification relies on extracting and utilizing shared visual features, i.e., shape, texture, and color, across categories. These fine-grained semantic details are crucial for distinguishing objects from the background, particularly as new categories appear. Research [31, 40, 2] in transfer learning and domain adaptation highlights that effective generalization depends on the ability to leverage these common attributes. However, in IOD, the evolving background class with each new task complicates the traditional models struggle to consistently capture and apply these shared features.
@@
Recent advances in vision-language models [32, 59, 58, 46, 43, 50, 51, 48, 49] offer promising solutions to these challenges. By integrating visual and textual data, these models provide a richer contextual understanding of both objects and their backgrounds. This cross-modal approach enhances the ability to retain and align shared semantic information across tasks, thereby mitigating the effects of background drift and improving the robustness of incremental object detection. The inability to maintain a coherent representation of these semantic relationships across tasks is thus a fundamental problem in IOD, compromising its overall performance as it encounters new categories.
@@
In response to these challenges, we propose a novel approach that leverages vision-language foundation models to address the background drift problem in IOD. Our method named Class-Agnostic Shared Attributes (CASA), a framework designed to capture and utilize common semantic information across incremental classes. By employing large language models, we generate candidate textual attributes relevant to the object categories in each detection task, curate these attributes based on their relevance to the current training data, and record their importance in an attribute assignment matrix. For subsequent tasks, we freeze the retained attributes while continuing to select and update relevant attributes, ensuring that the model adapts incrementally without losing previously learned knowledge. Moreover, we utilize OWL-ViT [30] as the baseline, preserving the original parameters of the pre-trained foundation model. Our method achieves only a minimal increase in parameter storage (0.7%) through parameter-efficient fine-tuning, significantly enhancing the scalability and adaptability of IOD. Extensive experiments on the COCO dataset demonstrate the effectiveness of our approach, achieving state-of-the-art performance across both two-phase and multi-phase incremental learning scenarios.
@@
  • We propose the Class-Agnostic Shared Attributes (CASA) for leveraging common semantic information across categories in incremental object detection, enhancing the ability to handle background drift.
@@
  • Extensive two-phase and multi-phase experiments on the COCO dataset demonstrate the effectiveness and efficiency of our approach, achieving state-of-the-art performance in incremental object detection.
2 Related Works
2.1 Vision-Language Models
@@
Vision-language models have emerged as powerful tools for understanding and integrating visual and textual data, enabling more comprehensive and context-aware representations of objects and scenes. One of the most prominent models is CLIP (Contrastive Language-Image Pretraining) [32], which leverages large-scale image-text pairs to learn a joint embedding space where visual and textual modalities are aligned. This approach has inspired subsequent research in vision-language models, such as ALIGN [14] and Florence [47], which further enhance the ability to generalize across diverse datasets and tasks. Recent advancements in vision-language models about object detection and classification have significantly improved robustness and adaptability, particularly in scenarios involving novel or evolving categories [32, 58, 60]. Among these methods, OWL-ViT [30], an open-world learning vision transformer, stands out by enabling more flexible and scalable detection and classification. FOMO [61] based on OWL-ViT utilize the attributes of known objects to recognize unknown objects in the open world environment.
2.2 Incremental Learning Based on Pre-trained Models
@@
Traditional incremental learning methods [6, 19, 21] usually start from a model trained from scratch. However, with the emergence of a variety of foundation models, leveraging their strong generalization to downstream tasks, incremental learning with pre-trained models has become an increasingly popular topic. It can be divided into three strategies: Prompt-based, Representation-based, and Model Mixture-based methods [56]. Prompt-based methods leverage the strong generalization capabilities of pre-trained models by using prompts to perform lightweight updates without fully fine-tuning all model parameters [42, 44, 45, 38, 15, 39, 16]. Representation-based methods directly utilize the generalization capabilities of PTMs to construct classifiers without making significant adjustments to the model itself [57, 28, 1, 52, 54, 12]. Model Mixture-based methods design a set of models during the learning process and employ techniques like model merging and ensemble learning to make final predictions [55, 41, 27]. By combining different models, these methods aim to capitalize on the strengths of each model to enhance overall learning performance.
2.3 Incremental Object Detection
@@
Incremental Object Detection (IOD) [37, 9] presents unique challenges compared to standard object detection, particularly in managing the evolving nature of the background class and ensuring that newly introduced categories do not interfere with previously learned ones. Existing incremental object detection methods can be broadly classified into three categories [63]: Knowledge Distillation (e.g., LWF [20]), Replay(e.g., iCaRL [33], ABR [24], and Regularization methods(e.g., ERD [9]). In recent years, with the popularity of the Transformer architecture, more and more methods try to use DETR as the baseline. CL-DETR [25] extends the DETR architecture to support incremental learning scenarios by leveraging transformer-based representations. Furthermore, CIOD [17] specifically focuses on maintaining detection accuracy across a growing number of classes by integrating adaptive feature extractors and regularization strategies. Additionally, the integration of vision-language models into IOD [18, 7] is a promising direction. These methods demonstrate that textual prompts can provide more semantic information for the representation of visual modality.
@@
Figure 1: Illustration of our proposed Class-Agnostic Shared Attribute (CASA) in Vision-Language Models for Efficient Incremental Object Detection. We leverage large language models to generate candidate textual attributes $E_{a}$ and then select the most relevant ones $\hat{E}_{a}^{t}$ based on the current training data, documenting their significance in an attribute assignment matrix $A^{t}$. In subsequent tasks, we retain and freeze these selected attributes, continuing the process by choosing from the remaining candidates and appending them after $\hat{E}_{a}^{t-1}$, and updating the attribute assignment matrix accordingly.
3 Method
3.1 Problem Formulation and Overview
Incremental object detection is an extension of object detection that allows for learning new categories without forgetting old categories. Specifically, when the model learns new classes during Task $T_{t}$, it must also maintain the ability to recognize the old classes ranging from Task $T_{1}$ to Task $T_{t-1}$. However, the challenge of background shift happens when objects from previous or future tasks in IOD are not labeled in the current task and are mistakenly categorized as part of the background class. Therefore, we propose an innovative method that leverages shared attributes within vision-language foundation models to enhance incremental object detection shown in Figure 1.
3.2 Attribute and Assignment Matrix Generation
Before encoding the text information, we first generate a large amount of attribute information [29, 62] corresponding to different object categories using the large language model (i.e., Gpt-3.5). Then we apply the attribute prompt template to obtain $N$ pieces of attribute text information, such as “object which (is/has/etc) Shape is squared.” Subsequently, we encode these texts using a Text Encoder [32] to obtain the attribute embedding, denoted as $E_{a}=[\mathbf{e}^{t}_{A_{1}},\mathbf{e}^{t}_{A_{2}},...,\mathbf{e}^{t}_{A_{N}}]\in\mathbb{R}^{D\times N}$. The attribute embedding $E_{a}$ is shared by current task and future tasks. We also use the vision encoder during each task to encode image information [8]. In our vision-language model, it is used together with $E_{a}$ for our efficient incremental object detection.
It is worth mentioning that not all attribute embedding is used in the detection process, and the selected attribute embedding has specific weights for each class. Therefore, assigning a score to represent the attribute assignment for each class during the selection becomes particularly important. Therefore, we introduce an assignment matrix, where each element represents the score of an attribute embedding assigned in the detection process for a certain class. In the process of incremental object detection, we denote the assignment matrix in $T_{t}$ as $A^{t}\in\mathbb{R}^{N\times C^{1:t}}$, where $C^{1:t}$ represents the number of all categories learned until Task $T_{t}$, while $C^{t}$ represents the number of new categories in Task $T_{t}$. In each task, We first initialize this assignment matrix with arbitrary values between 0 and 1 for subsequent training.
3.3 Assignment Matrix Updating
In Task $T_{t}$, with attribute embedding $\mathbf{e}^{t}_{A}$ and visual embedding $\mathbf{e}^{v}$ [8], we can calculate the similarity vector $S^{t}$ as follows:
@@ 3.3 Assignment Matrix Updating
Subsequently, we can use the similarity vector $S^{t}$ to update the assignment matrix $A^{t}$. We take the newly introduced last $C^{t}$ columns of the matrix $A^{t}$ as $A^{t-1:t}$. Then, the class probability $p_{cls}$ corresponding to $C^{t}$ classes introduced in Task $T_{t}$ can be computed using the following equation [23]:
@@
Next, we can calculate the Binary Cross-Entropy (BCE) loss between the logits of these categories and the one-hot encoded targets, $y$, to update the assignment matrix $A^{t-1:t}$. Additionally, we should add an regularization loss to enforce sparsity, resulting in a more sparse matrix $A^{t-1:t}$. The loss function $\mathcal{L}_{upd}$ can be formulated as:
$\mathcal{L}_{upd}=\mathcal{L}_{BCE}(p_{cls},y)+\lambda_{1}\sum_{i,j}|A^{t-1:t}_{t_{i,j}}|,$ (3)
where $\lambda_{1}$ is a tunable hyperparameter, which is set to 0.01 fixed in our work.
@@
At this stage, we have obtained an assignment matrix $A^{t-1:t}\in\mathbb{R}^{N\times C^{t}}$ with elements ranging between 0 and 1. Next, we proceed to filter the shared attributes. We assume $H_{a}$ represents the number of representative attributes selected for each category, consistently set to 25 in our experiments. We flatten $A^{t-1:t}$ into a one-dimensional vector with $N\times C^{t}$ elements, from which we select the top $C^{t}\times H_{a}$ elements with the highest values. The remaining elements are set to zero. Afterwards, we reshape $A^{t-1:t}$ back into a two-dimensional matrix where all elements are either 0 or 1.
@@
For the first task, we save the assignment matrix $A^{1}$. For the following tasks, to implement incremental learning, we concatenate the $A^{t-1}$ matrix, which is saved from the previous task, with the current task’s $A^{t-1:t}$ matrix row-wise to form a new $N\times C^{1:t}$ binary matrix $A^{t}$. In Task $T_{t}$, it is essential that the assignment matrix $A^{t}$ matrix is saved for following incremental object detection.
3.4 Attribute and Assignment Matrix Selection
At this point, in Task $T_{t}$ each value in a row of the assignment matrix $A^{t}$ represents whether a particular attribute is meaningful for detecting a specific category. If the value in a row corresponding to a category is 1, it signifies that this attribute is relevant for detecting that category. Conversely, the value of 0 indicates irrelevance. We then mask the attributes according to $A^{t}$ that all values in the row of the assignment matrix $A^{t}$ are 0.
We iterate through each row of the $A^{t}$ matrix and find $\hat{N}^{t}$ rows where the elements are not all zeros. In principle, we would select $C^{1:t}\times H_{a}$ rows attribute embedding $E_{a}$ in Task $T_{t}$, but actually how much we select is far fewer than $C^{1:t}\times H_{a}$. This is because some attributes are shared among classes within Task $T_{t}$. More importantly, as incremental learning progresses, an increasing number of attributes are shared between Task $T_{t}$ and all previous tasks. We need to retain these rows and delete the rows that are all zeros. Thus, we look for the indices of the rows where the element values are not all zeros, denoted as $id^{raw}_{t}$. To achieve incremental object detection, we refer to the indices $id_{t-1}$ saved from the previous task. We then check whether each index in $id^{raw}_{t}$ appears in $id_{t-1}$. The indices that do not appear in $id_{t-1}$ are added to the end of $id_{t-1}$, forming the indices $id_{t}$ for the current task. Therefore, to achieve incremental object detection for Task $T_{t}$, we need to save $id_{t-1}$ in Task $T_{t-1}$.
Next, we use $id_{t}$ to filter and select the attribute embedding $E_{a}$ into $\hat{E}_{a}^{t}$ and assignment matrix for Task $T_{t}$, retaining only the rows corresponding to the indices recorded in $id_{t}$ to form the $\hat{E}_{a}^{t}$ and $A^{t}$. During this filtering process, we may find that some indices in $id_{t}$ also appear in $id_{t-1}$, which indicates that the attributes selected for the current task have already been selected in previous tasks. This achieves attribute sharing under class-agnostic conditions.
3.5 Adaption for Attribute Refinement
After this step, we refine the attribute embedding $\hat{E}_{a}^{t}$ based on $A^{t-1:t}$ for the current task. However, the textual attribute information generated from the text and the visual information are orthogonal in space, belonging to different domains. Therefore, before fine-tuning the textual attribute information for visual information, we adapt the attribute embedding $\hat{E}_{a}^{t}$ to eliminate barriers between visual and textual information. For each category in the current task, we take $M$ samples and calculate the visual mean embedding $\bar{E}^{v}$ for these $M$ samples:
$\bar{E}^{v}(c)=\frac{1}{M}\sum_{i=0}^{M}\mathbf{e}^{v}_{i}.$ (4)
$+\lambda_{2}\mathcal{L}_{MSE}\left(\hat{E}_{a}^{t}[\ :Q,:],\hat{E}_{a}^{t-1}\right),$
where $\lambda_{2}$ is also a tunable hyperparameter. In this way, we adapt the attribute embedding $\hat{E}_{a}^{t}$, eliminating barriers between visual and textual information, which can be used for future refinement.
3.6 Refinement for Inference
In Task $T_{t}$, after adapting the attribute information, we perform refining on a small scale to better apply the assignment matrix $A^{t}$ and the attribute embedding $\hat{E}_{a}^{t}$ to the subsequent inference. Similar to Equation 1, we first compute the similarity vector $\hat{S}_{t}$ between the attribute embedding $\hat{E}_{a}^{t}$ and visual embedding $\mathbf{e}^{v}$ in the current task. Next, with the assignment matrix $A^{t-1:t}$, logits corresponding to these $C_{t}$ categories can be computed using Equation 2. Unlike the previous process, at this stage, we need to keep the assignment matrix $A^{t}$ unchanged and update $\hat{E}_{a}^{t}$. We calculate the BCE loss between the probabilities $P$ of these categories and the targets $T$, which are a list of labels corresponding to the target categories in the image.
@@ 3.6 Refinement for Inference
Table 1: CASA results (%) on COCO 2017 in two-phase setting 70+10. The best performance is highlighted in bold.
@@
$\mathcal{L}_{ref}=$ $\ \mathcal{L}_{BCE}\left(P,U\right)+\lambda_{3}\mathcal{L}_{MSE}\left(\hat{E}_{a}^{t}[\ :Q,:],\hat{E}_{a}^{t-1}\right),$ (6)
where $\lambda_{3}$ is also a tunable hyperparameter. After refining the attribute embedding, the $\hat{E}_{a}^{t}$ and the $A^{t}$ can be used in the following inference stage.
In the inference phase, similar to the process described in Equation 1 and Equation 2, the image to be inferred is first encoded by the Vision Transformer Encoder to obtain visual embedding $\mathbf{e}^{v}$. The encoded vectors $\mathbf{e}^{v}$ are then multiplied by the $\hat{E}_{a}^{t}$ obtained during model training, resulting in the similarity vectors. Subsequently, they are multiplied by the assignment matrix $A^{t}$ to obtain the probability of the specific category.
4 Experiments
4.1 Dataset and Evaluation Metrics
We evaluate our method on the MS COCO 2017 dataset, which is widely used in recent incremental object detection work. Via various experimental settings, both in two-phase and multi-phase, we conduct a comprehensive comparison with other incremental object detection methods in terms of evaluation metrics $AP$, $AP_{.5}$ and $AP_{.75}$, demonstrating the superiority of our method. We also evaluate the metric called Forgetting Percentage Points ($FPP$) followed by CL-DETR [25]: $FPP=AP^{1}-AP^{t}_{old}$, where $AP^{1}$ evaluates for all classes in the first task and $AP^{t}_{old}$ evaluates for previous classes learned from the first task during current task. Moreover, through the comparison of False Positives ($FP$), we show that our method CASA effectively overcomes the background shift problem [24] in incremental object detection.
4.2 Implementation Details

### v3

@@ Abstract
Incremental object detection is fundamentally challenged by catastrophic forgetting. A major factor contributing to this issue is background shift, where background categories in sequential tasks may overlap with either previously learned or future unseen classes. To address this, we propose a novel method called Class-Agnostic Shared Attribute Base (CASA) that encourages the model to learn category-agnostic attributes shared across incremental classes. Our approach leverages an LLM to generate candidate textual attributes, selects the most relevant ones based on the current training data, and records their importance in an assignment matrix. For subsequent tasks, the retained attributes are frozen, and new attributes are selected from the remaining candidates, ensuring both knowledge retention and adaptability. Extensive experiments on the COCO dataset demonstrate the state-of-the-art performance of our method. footnotetext: Supported by Shenzhen Peacock Team KQTD 20240729102051063 and the China Postdoctoral Science Foundation under Grant Number BX20240013 and 2024M760113. Equal Contribution: * ; Corresponding author: †
Index Terms:
incremental object detection learning, vision-language models, efficient learning
I Introduction
@@
Object detection models have achieved remarkable success in recognizing and localizing objects within a fixed set of categories. They assume a static training environment where all object categories are known and labeled beforehand. In real-world scenarios, as new object categories emerge over time, the need for continuous updates to these models becomes crucial—a process known as incremental object detection (IOD) [1, 2]. However, IOD also suffers significant challenges, particularly in managing background drift [3]. Background drift occurs when objects belonging to previous or future tasks in IOD are not annotated in the current task and are instead assigned to the background class. This can cause serious misclassification issues [2], as the model may later confuse these background objects with newly introduced categories or fail to recognize them altogether. This highlights a key challenge in IOD: how to maintain the ability to generalize across tasks without comprehensive annotations.
@@
In traditional object detection, accurate classification relies on extracting and utilizing shared visual features, i.e., shape, texture, and color. These fine-grained semantic details are crucial for distinguishing objects from the background, particularly as new categories appear. Research in transfer learning and domain adaptation highlights that effective generalization depends on the ability to leverage these common attributes. However, in IOD, the evolving background class with each new task complicates the traditional models struggle to consistently capture and apply these shared features.
@@
Recent advances in vision-language models [4, 5] offer promising solutions to these challenges. By integrating visual and textual data, these models provide a richer contextual understanding of both objects and their backgrounds. This cross-modal approach enhances the ability to retain and align shared semantic information across tasks, thereby mitigating the effects of background drift and improving the robustness of IOD. The inability to maintain a coherent representation of these semantic relationships across tasks is thus a fundamental problem in IOD, compromising its overall performance as it encounters new categories.
@@
To tackle these challenges, we propose Class-Agnostic Shared Attributes (CASA), a novel approach that leverages vision-language foundation models to address background drift. It captures and utilizes common semantic information across incremental classes. By employing LLM, we generate candidate textual attributes relevant to the object categories, curate these attributes based on their relevance to the current training data, and record their importance in an attribute assignment matrix. For subsequent tasks, we freeze the retained attributes while continuing to select and update relevant attributes, ensuring that the model adapts incrementally without losing previously learned knowledge. Building on OWL-ViT [6], CASA achieves only a minimal increase in parameter storage (0.7%) through parameter-efficient fine-tuning, significantly enhancing its scalability and adaptability. Experiments on the COCO dataset demonstrate its effectiveness, achieving state-of-the-art performance across both two-phase and multi-phase incremental learning scenarios.
@@
  • We propose Class-Agnostic Shared Attributes (CASA) for leveraging common semantic information across categories in IOD, overcome the background drift.
@@
  • Extensive two-phase and multi-phase experiments on COCO dataset demonstrate the effectiveness and efficiency of CASA, achieving SOTA performance in IOD.
II Related Works
II-A Vision-Language Models
@@
Vision-language models have emerged as powerful tools for understanding and integrating visual and textual data, enabling more comprehensive and context-aware representations of objects and scenes. One of the most prominent models is CLIP (Contrastive Language-Image Pretraining) [4], which leverages large-scale image-text pairs to learn a joint embedding space where visual and textual modalities are aligned. This approach has inspired subsequent research in vision-language models, such as ALIGN and Florence, which further enhance the ability to generalize across diverse datasets and tasks. Recent advancements in vision-language models about object detection and classification have significantly improved robustness and adaptability, particularly in scenarios involving novel or evolving categories [4]. Among these methods, OWL-ViT [6], an open-world learning vision transformer, stands out by enabling more flexible and scalable detection and classification. FOMO [7] based on OWL-ViT utilizes the attributes of known objects to recognize unknown objects in the open-world environment.
II-B Incremental Learning with Pre-trained Models (PTMs)
@@
Traditional incremental learning methods [8, 9, 10, 11] usually start with a model trained from scratch. With the emergence of a variety of foundation models, incremental learning with PTMs aims to leverag their strong generalization to downstream tasks. It can be divided into three strategies: Prompt-based, Representation-based, and Model Mixture-based methods [12]. Prompt-based methods leverage the strong generalization capabilities of PTMs by using prompts to perform lightweight updates without fully fine-tuning all model parameters [13]. Representation-based methods directly utilize the generalization capabilities of PTMs to construct classifiers without making significant adjustments to the model itself [14]. Model Mixture-based methods design a set of models during the learning process and employ techniques like model merging and ensemble learning to make final predictions [15, 16]. By combining different models, these methods aim to capitalize on the strengths of each model to enhance overall learning performance.
II-C Incremental Object Detection
@@
Incremental Object Detection [17, 18, 9, 19, 20] presents unique challenges compared to standard object detection, particularly in managing the evolving nature of the background class and ensuring that newly introduced categories do not interfere with previously learned ones. Existing IOD methods can be broadly classified into three categories: Knowledge Distillation (e.g., LWF [18]), Replay(e.g., iCaRL [9], ABR [2]) and Regularization methods(e.g., ERD [1], LID [21]). In recent years, with the popularity of the Transformer architecture, more and more methods have tried to use DETR as the baseline. CL-DETR [19] extends the DETR architecture to support incremental learning scenarios by leveraging transformer-based representations. Furthermore, CIOD [20] specifically focuses on maintaining detection accuracy across a growing number of classes by integrating adaptive feature extractors and regularization strategies. Additionally, the integration of vision-language models into IOD is a promising direction. These methods demonstrate that textual prompts can provide more semantic information for the representation of visual modality.
@@
Figure 1: Illustration of our proposed Class-Agnostic Shared Attribute (CASA). We leverage LLMs to generate the shared attribute base $E_{a}$ and then select the most relevant ones $\hat{E}_{a}^{t}$ based on the current training data, documenting their significance in an attribute assignment matrix $A^{t}$. In subsequent tasks, we retain and freeze these selected attributes, continuing the process by choosing from the remaining candidates and appending them after $\hat{E}_{a}^{t-1}$, and updating the attribute assignment matrix.
III Method
III-A Problem Formulation and Overview
IOD is an extension topic of conventional object detection, allowing for learning new categories without forgetting old ones. Specifically, when the model learns a new class during Task $T_{t}$, it should retain its ability to recognize the old classes ranging from Task $T_{0}$ to Task $T_{t-1}$. In this work, we propose encouraging the model to learn category-agnostic shared attributes across different objects, rather than overfitting to class-specific features, which can result in catastrophic forgetting. The pipeline of our method is shown in Figure 1.
III-B Shared Attributes Base Creation
In this module, we aim to generate the class-agnostic attributes base using the data from the first-given task ($T_{0}$). Specifically, we first generate a large amount of attributes corresponding to different objects using LLMs (i.e., Gpt-4o). Then we apply the attribute prompt template to obtain $N$ pieces of attribute texts, such as “object which has color is red.” Note that the used prompts are designed to be class-agnostic, where we use abstract “object” to replace the specific name of the object. This manner encourages the model to learn more common and shared attributes across different objects. Subsequently, we compress these texts with a Text Encoder [4] to obtain the attribute embedding base, denoted as $E_{a}=[\mathbf{e}^{t}_{A_{1}},\mathbf{e}^{t}_{A_{2}},...,\mathbf{e}^{t}_{A_{N}}]\in\mathbb{R}^{D\times N}$. The created attribute embedding base $E_{a}$ will be frozen and used for all tasks. Note that not all attributes are informative and general for our task.
III-C Training
III-C1 Motivation
The primary challenge in IOD is background shift, where the model tends to overfit to object-specific visual features from previous tasks, limiting its ability to retain prior knowledge and adapt to new, unseen classes.
To address this challenge, we propose leveraging vision-language models to learn a class-agnostic shared attributes base, which can be utilized for more general and robust detection. This approach involves solving two core problems: (a) How can we identify the most informative and general attributes from the shared attribute base? (b) How can we effectively utilize these selected attributes for object detection? The following sections provide detailed technical methodologies.
III-C2 Attributes Selection and Attributes Assignment
To address the first problem, we identify the most informative and general attributes by matching attributes embeddings with the visual embeddings ($\mathbf{e}^{v}$) extracted from the visual encoder. Specifically, during Task $T_{t}$, given the attributes embedding $\mathbf{e}^{t}_{A}$ and the visual embedding $\mathbf{e}^{v}$, the similarity vector $S^{t}$ is computed as:
@@ III-C2 Attributes Selection and Attributes Assignment
Here, the similarity vector $S^{t}$ reflects the relevance of each attribute, with higher values indicating more informative attributes. This step ensures the selection of attributes that are meaningful for object detection.
To map the selected attributes to object categories for detection, we introduce an assignment matrix, where each element of the assignment matrix represents the score of an attribute being associated with a specific category. In Task $T_{t}$, the assignment matrix is defined as $A^{t}\in\mathbb{R}^{N\times C^{1:t}}$, where $C^{1:t}$ denotes the total number of categories learned up to Task $T_{t}$, and $C^{t}$ represents the number of new categories introduced in Task $T_{t}$. Initially, the assignment matrix is initialized with arbitrary values between 0 and 1.
The similarity vector $S^{t}$ is then used to update the assignment matrix $A^{t}$. Specifically, for the newly introduced classes in Task $T_{t}$, we extract the last $C^{t}$ columns of the assignment matrix, denoted as $A^{t-1:t}$. The class probability $p_{cls}$ for these new classes is computed as:
@@
Subsequently, the Binary Cross-Entropy (BCE) loss is calculated between the predicted probabilities $p_{cls}$ and the one-hot encoded targets $y$. Additionally, we add a regularization loss to enforce sparsity. The overall loss function $\mathcal{L}_{upd}$ can be formulated as:
$\mathcal{L}_{upd}=\mathcal{L}_{BCE}(p_{cls},y).)+\lambda\sum_{i,j}|A^{t-1:t}_{t_{i,j}}|,$ (3)
where $\lambda$ is a tunable hyperparameter, which is set to 0.01 fixed in our work. This loss is used to iteratively refine the assignment matrix $A^{t-1:t}$, ensuring that it captures meaningful relationships between attributes and the new categories.
III-C3 Attribute Filtering and Incremental Learning
@@
Once the assignment matrix $A^{t-1:t}$ has been updated, we proceed to filter the shared attributes. Specifically, for each category, we select the top $H_{a}$ representative attributes, where $H_{a}$ is a predefined constant. To achieve this, $A^{t-1:t}$ is flattened into a one-dimensional vector, and the top $C^{t}\times H_{a}$ elements with the highest scores are selected. The remaining elements are set to zero, and the matrix is reshaped back into a binary matrix, where each element is either 0 or 1.
@@
To implement incremental learning, the assignment matrix from the previous task, $A^{t-1}$, is concatenated with the current matrix $A^{t-1:t}$ to form a new matrix $A^{t}\in\mathbb{R}^{N\times C^{1:t}}$. This updated matrix is saved for subsequent tasks, enabling seamless incremental learning across multiple tasks.
@@
TABLE I: CASA results (%) on COCO 2017 in two-phase setting 70+10. The best performance is highlighted in bold.
@@
III-C4 Attribute Sharing and Class-Agnostic Learning
During IOD, the assignment matrix $A^{t}$ allows for attribute sharing across tasks. For any row in $A^{t}$, a value of 1 indicates that the corresponding attribute is relevant for a specific category, while a value of 0 indicates irrelevance. Rows in $A^{t}$ where all values are 0 are removed, ensuring that only useful attributes are retained. Additionally, attributes shared across tasks are preserved by maintaining indices from the previous task, denoted as $id_{t-1}$. These indices are updated to form $id_{t}$ for the current task. Using $id_{t}$, we filter the attribute embeddings and assignment matrix, retaining only rows corresponding to meaningful indices. This process ensures that shared attributes are utilized effectively, enabling the model to learn in a category-agnostic manner and improving its generalization across tasks.
Besides, there are another two steps written in supplementary materials II. By addressing both attribute selection and utilization, our method effectively mitigates the challenge of background shift and enables robust IOD.
III-D Inference
During inference, the trained model utilizes the category-agnostic shared attributes and the assignment matrix to detect objects incrementally across tasks. Given an input image, the visual encoder extracts the visual embeddings $\mathbf{e}^{v}$, which are then matched with the attribute embeddings $\mathbf{e}_{A}^{t}$ using the similarity computation defined in Equation 1. This results in a similarity vector $S^{t}$, which highlights the relevance of each attribute for the given input. Using the assignment matrix $A^{t}$ from the current task, the model maps the similarity vector $S^{t}$ to the class probabilities $p_{cls}$ for all categories learned up to the current task $T_{t}$. This mapping is performed as described in Equation 2, ensuring that the output reflects the contributions of both new and previously learned attributes. The resulting probabilities are used to classify the detected objects into one of the learned categories, while maintaining consistency with the knowledge retained from earlier tasks.
By leveraging the binary structure of the assignment matrix $A^{t}$, the model dynamically filters irrelevant attributes, focusing only on those that are meaningful for object detection. Additionally, the preserved shared attributes across tasks ensure that the model generalizes well to new categories without forgetting prior knowledge. This process enables the detection of objects from both old and new categories in a seamless and efficient manner, demonstrating the effectiveness of the proposed method in addressing the challenges of IOD.
IV Experiments
IV-A Dataset and Evaluation Metrics
We evaluate our method on the MS COCO 2017 dataset, which is widely used in IOD. Via both two-phase and multi-phase setting, we conduct a comprehensive comparison with other IOD methods in terms of evaluation metrics $AP$, $AP_{.5}$ and $AP_{.75}$. We also evaluate the metric called Forgetting Percentage Points ($FPP$) followed by CL-DETR [19]: $FPP=AP^{1}-AP^{t}_{old}$, where $AP^{1}$ evaluates for all classes in the first task and $AP^{t}_{old}$ evaluates for previous classes learned from the first task during current task. The comparison of False Positives ($FP$) shows that CASA effectively overcomes the background shift problem [2] in IOD.
IV-B Implementation Details